# To curve, or not to curve: Is curvature-assisted quintessence observationally viable?

### v2

@@ 4.2 Bayesian analysis and parameter estimation
We now perform a full Bayesian statistical analysis of curvature-assisted, single-field quintessence to scan the entire parameter space of the model and constrain, particularly, the $\lambda$ parameter. This will tell us whether large values of $\lambda$ (i.e., $\lambda>\sqrt{2}$) preferred by the statement (2) become observationally viable when the spatial curvature of the Universe is nonzero. We have already argued, based on qualitative discussions, as well as through a numerical grid analysis of the phase space of the model, that it is not the case, but here, we demonstrate that through a detailed comparison of the predictions of the model with some of the most recent cosmological observations. We employ the cosmic microwave background (CMB) distance priors provided by Planck Efstathiou and Bond (1999); Nesseris and Perivolaropoulos (2007); Aghanim et al. (2020), a baryon acoustic oscillations (BAO) data compilation Alestas et al. (2022), and the Pantheon$+$ Type Ia supernovae (SNe Ia) data Riess et al. (2022). [^2] We do not use the recent constraints provided by the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024), in order to avoid potential systematic effects, as the official DESI likelihood is not yet publicly available. In any case, pre-DESI cosmological data are sufficient for the purpose of the present paper, which is to show that large values of the parameter $\lambda$ are excluded observationally even if the Universe is negatively curved—the DESI results would not increase the upper bound on $\lambda$ that we report here. The distance prior parameters are comprised of the acoustic scale $I_{\mathrm{A}}$, which measures the temperature of the CMB in the transverse direction, the shift parameter $R$, which measures the peak spacing of the temperature in the power spectrum, and the baryon density $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}$, where $\Omega_{\mathrm{b}}$ is the baryon density parameter and $h\equiv H_{0}/100$ with $H_{0}$ the present value of the Hubble expansion rate $H$. As for the BAO compilation, we use the data from WiggleZ Blake et al. (2012); Escamilla-Rivera (2016), 6dFGS Beutler et al. (2011), the Dark Energy Survey (DES) Abbott et al. (2022), and the fourth generation of the Sloan Digital Sky Survey (SDSS-IV) Alam et al. (2021); du Mas des Bourboux et al. (2020).
@@
We perform a Markov chain Monte Carlo (MCMC) scan of the parameter space of the quintessence model for two cases. [^3] The Mathematica codes used in our analysis will be made publicly available in a GitHub repository upon publication of the paper. We first consider a flat universe, i.e., we fix the present value of the curvature density parameter $\Omega_{k,0}$ to zero. We then allow $\Omega_{k,0}$ to vary as a free parameter. For each case, we perform two separate analyses, one by setting the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ to zero and one by varying it as a free parameter. In both cases, we fix $\phi_{0}$ (the initial value of the scalar field) to zero and vary the parameter $V_{0}$ in the potential $V(\phi)=V_{0}e^{-\lambda\phi}$, as one of the two parameters is redundant and can be replaced by the other. We impose flat (or uniform) priors on all free parameters—table 1 shows the ranges of parameters we employ in our scans. Comparing the parameter constraints we obtain for different cases will tell us whether assisting the accelerated expansion with a nonzero spatial curvature will increase the upper bound on the parameter $\lambda$ so that large values suggested by the statement (2) become observationally viable.
@@
Figure 5: The $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions for the quintessence model of dark energy with an exponential potential when the spatial curvature of the Universe is set to zero (i.e., $\Omega_{k,0}=0$, where $\Omega_{k,0}$ is the present value of the curvature density parameter $\Omega_{k}$). We show the results for two cases: 1) the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ is set to zero, and 2) $\dot{\phi}_{0}$ is varied as a free parameter.
Figure 6: As in fig. 5, but for cases where the spatial curvature of the Universe is allowed to be nonzero, i.e., the present value of the curvature density parameter, $\Omega_{k,0}$, is varied as a free parameter.
@@
We present the results of our MCMC analysis in figs. 5 and 6, where two-dimensional $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions (PDFs) for different parameters and different cases are provided. We also present in table 2 $1\sigma$ (i.e., $68.3\%$ confidence level) constraints on the parameters. Since the parameter of interest here is $\lambda$, let us compare our constraints on that parameter as given in the figures and the table for the different cases we have studied. Fig. 5 and table 2 show that cosmological observations place an upper bound of $\sim 0.6$ on $\lambda$ when $\dot{\phi}_{0}$ is fixed to zero, and an upper bound of $\sim 0.5$ when $\dot{\phi}_{0}$ is varied, at the $63.3\%$ confidence level, which are consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019). When $\Omega_{k,0}$ is allowed to vary, we obtain the upper bound of $\sim 0.7$ on $\lambda$ (as given in fig. 6 and table 2), independently of whether $\dot{\phi}_{0}$ is varied or fixed to zero, even though the contours and the one-dimensional PDFs in the two cases are slightly different. Note, however, that it is important to vary $\dot{\phi}_{0}$ in the case of free $\Omega_{k,0}$ for full exploration of the phase space of the quintessence model—there is no a priori reason for setting it to zero. It is, therefore, true that allowing $\Omega_{k,0}$ to vary pushes the upper bound on $\lambda$ to larger values, but the improvement is tiny and the favored values of $\lambda$ are far below the $\lambda=\sqrt{2}$ limit of the statement (2). This means that the new fixed point in the phase space of the model which arises as a result of $\Omega_{k}$ taking positive values, i.e., the Universe having a negative curvature, does not help with making $\lambda>\sqrt{2}$ consistent with observations—this is what we expected from the qualitative analyses of previous sections. It is, however, interesting to note that when $\Omega_{k,0}$ is allowed to vary, cosmological data seem to prefer nonzero curvature and therefore indicate a slight preference for an open universe, although the preference is only slightly more than $2\sigma$. We do not perform a model comparison here, but it is likely that the slightly larger upper bound on $\lambda$ for the curvature-assisted model compared to the flat one is due to the increase in the number of free parameters (i.e., the number of degrees of freedom) in the former case. As stated above, the results of the full MCMC analysis perfectly agree with the findings of the numerical grid analysis of section 4.1.
@@ 5 Conclusions
We have first explained the importance of scalar fields in string theory, and after reviewing the theoretical reasons for why asymptotic corners of moduli space of string theoretic EFTs are interesting, we have argued that neither a de Sitter vacuum nor a quintessence scenario with a nearly flat potential can be realized in those corners. The reason is that all known asymptotic string theory examples seem to imply that there is a lower bound on the gradient of the potential that is too steep to provide accelerated expansion consistent with cosmological observations—this has led to a conjecture that this steepness is a universal feature of quantum gravity in the asymptotic regions of moduli space, and therefore, one may need to focus instead on the bulk of the space. However, the authors of the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) have argued that one could obtain observationally viable accelerated expansion even for steep potentials, i.e., from a fully controlled string theoretic EFT, if the Universe is negatively curved.
We have focused on the single-field quintessence model of dark energy with the exponential potential $V(\phi)=V_{0}e^{-\lambda\phi}$, and after reviewing the phase space of the model and the corresponding dynamical system, we have shown, both qualitatively and through a Bayesian comparison of the model’s predictions with a set of recent cosmological data, that large values of $\lambda$, i.e., values larger than the lower bound $\sqrt{2}$ given by the strong de Sitter conjecture, are observationally excluded, even for nonzero values of the spatial curvature. We have demonstrated that through a numerical grid analysis of the phase space of the model, as well as a Bayesian analysis of the model where we have scanned the parameter space using an MCMC analysis and compared various cosmological quantities with their measured values.
@@
These results are consistent with our investigation of the phase space of the model (also allowing for an arbitrary value of the initial velocity of the scalar field), where we have performed a numerical grid analysis and shown that even though a new fixed point exists in the phase space when $\Omega_{k}>0$, none of the solutions of the set of differential equations describing the dynamical system simultaneously satisfy the two necessary conditions for a viable cosmology: 1) contain an extended period of matter domination before the onset of late-time cosmic acceleration, and 2) have an effective equation of state parameter that is $\sim-0.7$ at the present time.
@@
This all means that if the strong de Sitter conjecture is indeed true and a universal feature of the asymptotic regions of string theory, or more generally of any theories of quantum gravity, then there seem to be two avenues to explore if one intends to explain accelerated expansion in such theories: 1) construct accelerating solutions through the bulk of moduli space instead of the asymptotic corners of it, or 2) construct models with more than one scalar field, which have been shown to be able to provide observationally viable acceleration even with steep potentials (see, e.g., Refs. Achúcarro and Palma (2019); Cicoli et al. (2020a, b); Akrami et al. (2021); Eskilt et al. (2022)), although this too may be challenging in asymptotic corners of string theory due to universal features of the moduli space metric.
Note added: In the final stages of the present work, we became aware of Refs. Andriot et al. (2024); Bhattacharya et al. (2024), where similar results have been presented. In particular, the authors of Ref. Bhattacharya et al. (2024) have performed a cosmological data analysis of the curvature-assisted quintessence model where they have set $\dot{\phi}_{0}$ to zero. Although the results of our analysis show that the constraints on the parameters of the model do not strongly depend on whether $\dot{\phi}_{0}$ is fixed or is allowed to vary as a free parameter, this cannot be known a priori, as $\dot{\phi}_{0}$ is one of the phase space variables of the model and fixing it may result in missing important regions of the phase space when $\Omega_{k}$ takes nonzero values. We also became aware of Ref. Ramadan et al. (2024), where a cosmological analysis of single-field quintessence with an exponential potential has been performed for the zero-curvature model. In both cases, our results agree with theirs where they overlap.
@@ Acknowledgements.
We thank M. Raveri for helpful discussions. We also thank the pizza restaurant “Internacional cuisine” from the south of Madrid that provided us with delicious pizzas during the ChatCQG meetings that led to this work. G.A. is supported by the Spanish Attraccion de Talento contract 2019-T1/TIC-13177 granted by the Comunidad de Madrid. M.D. is supported by the grant FPI SEV-2016-0597-19-3 from the Spanish National Research Agency from the Ministry of Science and Innovation. I.R. is supported by the Spanish FPI grant PRE2020-094163. Y.A. acknowledges support by the Spanish Research Agency (Agencia Estatal de Investigación)’s grant RYC2020-030193-I/AEI/10.13039/501100011033, and by the European Social Fund (Fondo Social Europeo) through the Ramón y Cajal program within the State Plan for Scientific and Technical Research and Innovation (Plan Estatal de Investigación Científica y Técnica y de Innovación) 2017-2020. S.N. acknowledges support from the research project PID2021-123012NB-C43. M.M. is supported by an Atraccion del Talento Fellowship 2022-T1/TIC-23956 from Comunidad de Madrid. M.M. thanks the KITP program “What is String Theory?” for providing a stimulating enviroment for discussion. This research is supported in part by the grant NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP). M.D., I.R., and M.M. acknowledge the hospitality of the Department of Theoretical Physics at CERN and the Department of Physics of Harvard University during the different stages of this work. M.D. thanks the “Geometry, Strings and the Swampland Program” workshop organized by the Max Planck Institute for Physics and the Arnold Sommerfeld Center for Theoretical Physics for providing a fantastic environment for discussions during the early stages of this work. The authors thank the Spanish Research Agency (Agencia Estatal de Investigación) through the grants IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. The authors also acknowledge the use of the IFT Hydra cluster.

### v3

@@ 4.2 Bayesian analysis and parameter estimation
We now perform a full Bayesian statistical analysis of curvature-assisted, single-field quintessence to scan the entire parameter space of the model and constrain, particularly, the $\lambda$ parameter. This will tell us whether large values of $\lambda$ (i.e., $\lambda>\sqrt{2}$) preferred by the statement (2) become observationally viable when the spatial curvature of the Universe is nonzero. We have already argued, based on qualitative discussions, as well as through a numerical grid analysis of the phase space of the model, that it is not the case, but here, we demonstrate that through a detailed comparison of the predictions of the model with some of the most recent cosmological observations. We employ the cosmic microwave background (CMB) distance priors provided by Planck Efstathiou and Bond (1999); Nesseris and Perivolaropoulos (2007); Aghanim et al. (2020), a baryon acoustic oscillation (BAO) data compilation Alestas et al. (2022), and the Pantheon$+$ type Ia supernovae (SNe Ia) data Riess et al. (2022). [^2] We do not use the recent constraints provided by the Dark Energy Spectroscopic Instrument (DESI) Adame et al. (2024), in order to avoid potential systematic effects, as the official DESI likelihood is not yet publicly available. In any case, pre-DESI cosmological data are sufficient for the purpose of the present paper, which is to show that large values of the parameter $\lambda$ are excluded observationally even if the Universe is negatively curved—the DESI results would not increase the upper bound on $\lambda$ that we report here. The distance prior parameters are comprised of the acoustic scale $I_{\mathrm{A}}$, which measures the temperature of the CMB in the transverse direction, the shift parameter $R$, which measures the peak spacing of the temperature in the power spectrum, and the baryon density $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}$, where $\Omega_{\mathrm{b}}$ is the baryon density parameter and $h\equiv H_{0}/100$ with $H_{0}$ the present value of the Hubble expansion rate $H$. As for the BAO compilation, we use the data from WiggleZ Blake et al. (2012); Escamilla-Rivera (2016), 6dFGS Beutler et al. (2011), the Dark Energy Survey Abbott et al. (2022), and the fourth generation of the Sloan Digital Sky Survey Alam et al. (2021); du Mas des Bourboux et al. (2020).
@@
We perform a Markov chain Monte Carlo (MCMC) scan of the parameter space of the quintessence model for two cases. [^3] The Mathematica codes used in our analysis will be made publicly available in a GitHub repository upon publication of the paper. We first consider a flat universe; i.e., we fix the present value of the curvature density parameter $\Omega_{k,0}$ to zero. We then allow $\Omega_{k,0}$ to vary as a free parameter. For each case, we perform two separate analyses, one by setting the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ to zero and one by varying it as a free parameter. In both cases, we fix $\phi_{0}$ (the initial value of the scalar field) to zero and vary the parameter $V_{0}$ in the potential $V(\phi)=V_{0}e^{-\lambda\phi}$, as one of the two parameters is redundant and can be replaced by the other. We impose flat (or uniform) priors on all free parameters—Table 1 shows the ranges of parameters we employ in our scans. Comparing the parameter constraints we obtain for different cases will tell us whether assisting the accelerated expansion with a nonzero spatial curvature will increase the upper bound on the parameter $\lambda$ so that large values suggested by the statement (2) become observationally viable.
@@
Figure 5: The $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions for the quintessence model of dark energy with an exponential potential when the spatial curvature of the Universe is set to zero (i.e., $\Omega_{k,0}=0$, where $\Omega_{k,0}$ is the present value of the curvature density parameter $\Omega_{k}$). We show the results for two cases: (i) the initial value of the scalar field’s time derivative $\dot{\phi}_{0}$ is set to zero, and (ii) $\dot{\phi}_{0}$ is varied as a free parameter.
Figure 6: As in Fig. 5, but for cases where the spatial curvature of the Universe is allowed to be nonzero; i.e., the present value of the curvature density parameter, $\Omega_{k,0}$, is varied as a free parameter.
@@
We present the results of our MCMC analysis in Figs. 5 and 6, where two-dimensional $68.3\%$ and $95.5\%$ confidence regions and one-dimensional marginalized probability density functions (PDFs) for different parameters and different cases are provided. We also present in Table 2 $1\sigma$ (i.e., $68.3\%$ confidence level) constraints on the parameters. Since the parameter of interest here is $\lambda$, let us compare our constraints on that parameter as given in the figures and the table for the different cases we have studied. Figure 5 and Table 2 show that cosmological observations place an upper bound of $\sim 0.6$ on $\lambda$ when $\dot{\phi}_{0}$ is fixed to zero and an upper bound of $\sim 0.5$ when $\dot{\phi}_{0}$ is varied, at the $68.3\%$ confidence level, which are consistent with the results of Refs. Agrawal et al. (2018); Akrami et al. (2019); Raveri et al. (2019). When $\Omega_{k,0}$ is allowed to vary, we obtain the upper bound of $\sim 0.7$ on $\lambda$ (as given in Fig. 6 and Table 2), independently of whether $\dot{\phi}_{0}$ is varied or fixed to zero, even though the contours and the one-dimensional PDFs in the two cases are slightly different. Note, however, that it is important to vary $\dot{\phi}_{0}$ in the case of free $\Omega_{k,0}$ for full exploration of the phase space of the quintessence model—there is no a priori reason for setting it to zero. It is, therefore, true that allowing $\Omega_{k,0}$ to vary pushes the upper bound on $\lambda$ to larger values, but the improvement is tiny and the favored values of $\lambda$ are far below the $\lambda=\sqrt{2}$ limit of the statement (2). This means that the new fixed point in the phase space of the model which arises as a result of $\Omega_{k}$ taking positive values, i.e., the Universe having a negative curvature, does not help with making $\lambda>\sqrt{2}$ consistent with observations—this is what we expected from the qualitative analyses of previous sections. It is, however, interesting to note that when $\Omega_{k,0}$ is allowed to vary, cosmological data seem to prefer nonzero curvature and, therefore, indicate a slight preference for an open universe, although the preference is only slightly more than $2\sigma$. We do not perform a model comparison here, but it is likely that the slightly larger upper bound on $\lambda$ for the curvature-assisted model compared to the flat one is due to the increase in the number of free parameters (i.e., the number of degrees of freedom) in the former case. As stated above, the results of the full MCMC analysis perfectly agree with the findings of the numerical grid analysis in Section 4.1.
@@ 5 Conclusions
We have first explained the importance of scalar fields in string theory, and after reviewing the theoretical reasons for why asymptotic corners of moduli space of string theoretic EFTs are interesting, we have argued that neither a de Sitter vacuum nor a quintessence scenario with a nearly flat potential can be realized in those corners. The reason is that all known asymptotic string theory examples seem to imply that there is a lower bound on the gradient of the potential that is too steep to provide accelerated expansion consistent with cosmological observations—this has led to a conjecture that this steepness is a universal feature of quantum gravity in the asymptotic regions of moduli space, and, therefore, one may need to focus instead on the bulk of the space. However, the authors of the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) have argued that one could obtain observationally viable accelerated expansion even for steep potentials, i.e., from a fully controlled string theoretic EFT, if the Universe is negatively curved.
We have focused on the single-field quintessence model of dark energy with the exponential potential $V(\phi)=V_{0}e^{-\lambda\phi}$, and, after reviewing the phase space of the model and the corresponding dynamical system, we have shown, both qualitatively and through a Bayesian comparison of the model’s predictions with a set of recent cosmological data, that large values of $\lambda$, i.e., values larger than the lower bound $\sqrt{2}$ given by the strong de Sitter conjecture, are observationally excluded, even for nonzero values of the spatial curvature. We have demonstrated that through a numerical grid analysis of the phase space of the model, as well as a Bayesian analysis of the model where we have scanned the parameter space using an MCMC analysis and compared various cosmological quantities with their measured values.
@@
These results are consistent with our investigation of the phase space of the model (also allowing for an arbitrary value of the initial velocity of the scalar field), where we have performed a numerical grid analysis and shown that even though a new fixed point exists in the phase space when $\Omega_{k}>0$, none of the solutions of the set of differential equations describing the dynamical system simultaneously satisfy the two necessary conditions for a viable cosmology: (i) contain an extended period of matter domination before the onset of late-time cosmic acceleration and (ii) have an effective equation of state parameter that is $\sim-0.7$ at the present time.
@@
This all means that if the strong de Sitter conjecture is indeed true and a universal feature of the asymptotic regions of string theory, or more generally of any theories of quantum gravity, then there seem to be two avenues to explore if one intends to explain accelerated expansion in such theories: (i) construct accelerating solutions through the bulk of moduli space instead of the asymptotic corners of it, or (ii) construct models with more than one scalar field, which have been shown to be able to provide observationally viable acceleration even with steep potentials (see, e.g., Refs. Achúcarro and Palma (2019); Cicoli et al. (2020a, b); Akrami et al. (2021); Eskilt et al. (2022)), although this too may be challenging in asymptotic corners of string theory due to universal features of the moduli space metric.
Note added. In the final stages of the present work, we became aware of Refs. Andriot et al. (2024); Bhattacharya et al. (2024), where similar results have been presented. In particular, the authors of Ref. Bhattacharya et al. (2024) have performed a cosmological data analysis of the curvature-assisted quintessence model where they have set $\dot{\phi}_{0}$ to zero. Although the results of our analysis show that the constraints on the parameters of the model do not strongly depend on whether $\dot{\phi}_{0}$ is fixed or is allowed to vary as a free parameter, this cannot be known a priori, as $\dot{\phi}_{0}$ is one of the phase space variables of the model and fixing it may result in missing important regions of the phase space when $\Omega_{k}$ takes nonzero values. We also became aware of Ref. Ramadan et al. (2024), where a cosmological analysis of single-field quintessence with an exponential potential has been performed for the zero-curvature model. In both cases, our results agree with theirs where they overlap.
@@ Acknowledgements.
We thank M. Raveri for helpful discussions. We also thank the pizza restaurant “Internacional cuisine” from the south of Madrid that provided us with delicious pizzas during the ChatCQG meetings that led to this work. G.A. is supported by the Spanish Attraccion de Talento Contract No. 2019-T1/TIC-13177 granted by the Comunidad de Madrid. M.D. is supported by Grant No. FPI SEV-2016-0597-19-3 from the Spanish National Research Agency from the Ministry of Science and Innovation. I.R. is supported by the Spanish FPI Grant No. PRE2020-094163. Y.A. acknowledges support by the Spanish Research Agency (Agencia Estatal de Investigación)’s Grant No. RYC2020-030193-I/AEI/10.13039/501100011033, and by the European Social Fund (Fondo Social Europeo) through the Ramón y Cajal program within the State Plan for Scientific and Technical Research and Innovation (Plan Estatal de Investigación Científica y Técnica y de Innovación) 2017-2020. S.N. acknowledges support from the research Project No. PID2021-123012NB-C43. M.M. is supported by an Atraccion del Talento Fellowship 2022-T1/TIC-23956 from Comunidad de Madrid. M.M. thanks the KITP program “What is String Theory?” for providing a stimulating enviroment for discussion. This research is supported in part by the Grant No. NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP). M.D., I.R., and M.M. acknowledge the hospitality of the Department of Theoretical Physics at CERN and the Department of Physics of Harvard University during the different stages of this work. M.D. thanks the “Geometry, Strings and the Swampland Program” workshop organized by the Max Planck Institute for Physics and the Arnold Sommerfeld Center for Theoretical Physics for providing a fantastic environment for discussions during the early stages of this work. The authors thank the Spanish Research Agency (Agencia Estatal de Investigación) through the grants IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. The authors also acknowledge the use of the IFT Hydra cluster.